OpenStudy (anonymous):

two stonees are dropped from the top of a tower at half a second apart the time after dropping the first stone at which the distance between the two stones is 20m is??

OpenStudy (experimentx):

?? i did not understand your question ??? is it asking time or distance ??? also is distance between two stones 20m ??

OpenStudy (anonymous):

its about time??

OpenStudy (anonymous):

options are (a) 4s b 3.75s c 4.25s d 2s the answer is c can u solve it??

OpenStudy (experimentx):

\( s_1 - s_2 = 20 = (u_1t + 1/2 * at^2) - (u_2(t-1/2) + 1/2 * a(t-1/2)^2) \), u1, u2 = 0, a = g

OpenStudy (anonymous):

i did nt understand this +1/2∗a(t−1/2)2 can u clear it

OpenStudy (anonymous):

take g=10m/s2

OpenStudy (experimentx):

the time factor is decreased by half for second stone, i.e. T = t - 1/2

OpenStudy (amistre64):

how far does the first stone drop in half a second?

OpenStudy (anonymous):

how it came t^2 - (t-1/2)^2 = 4???

OpenStudy (experimentx):

well ... i simplified it greatly!!! .... use above relation!!!

OpenStudy (amistre64):

after half a second, the first stone can be thought of to have an initial velocity compared to that of the second stone. ..... p(t) = 9.8 t^2 v(t) = 4.9 t ; at t=1/2 = 2.45 p1(t) = 9.8 t^2 + 2.45t p2(t) = 9.8 t^2 when p1(t) - p2(t) = 20 we should be good

OpenStudy (amistre64):

plus half a second :)

OpenStudy (experimentx):

see you later ... !!!

OpenStudy (amistre64):

its best to interact in the post and not the messages of you have questions

OpenStudy (amistre64):

the position of the 1st stone under gravity can be modeled by the equation hmm, my gravity was off a little, but concept should be sound

OpenStudy (anonymous):

take g=10m/s2

OpenStudy (amistre64):

the equation for position is:\[p(t)=\frac{1}{2}g\ t^2+V_ot+H_o\] the derivative of position is velocity; \[v(t)=gt+V_o\]

OpenStudy (amistre64):

when the time = .5; the velocity of the first stone is at: 10+.5 = 5 right? so we adjust the equation and compare it to the second stone

OpenStudy (amistre64):

10*.5 = 5 cant type to save my life tonight

OpenStudy (amistre64):

\[p_1(t)=\frac{1}{2}g\ t^2+V_{o}t\] \[p_2(t)=\frac{1}{2}g\ t^2\] when the position between 1 and 2 = 20, we can solve for t \[p_1(t)-p_2(t)=20\] \[\frac{1}{2}g\ t^2+V_{o}t-\frac{1}{2}g\ t^2=20\] \[V_ot-20=0\] \[t=\frac{20}{V_o}\] total time elpsed is half a second more then

OpenStudy (amistre64):

20/5 = 4 + .5 = 4.5 but thats if you use g=10 instead of 9.8